# Non-ultralocal classical $r$-matrix structure

### v1

@@ 1 Introduction
Integrable 1+1 field theories are actively studied at the classical and quantum levels due to their deep relation to gauge theories, strings and conformal field theories, existence of soliton solutions and many other remarkable properties [16, 33, 25, 22, 23, 13, 15]. We hope to include into consideration a large class of models obtained as 1+1 field generalizations of the widely known finite-dimensional integrable systems of Calogero-Ruijsenaars family. In 1+1 field theories we deal with the fields depending on the time variable $t$ and the space variable $x$. It is assumed that all fields $\varphi(t,x)$ are periodic functions on a unit circle, i.e. $\varphi(t,x+2\pi)=\varphi(t,x)$. Integrable field analogues of the finite-dimensional many-body systems are known beginning from 2d Toda field theory [27]. In contrast to the Toda model the field generalization of the Calogero-Moser model is highly non-trivial (see also [38] for the field analogue of the Ruijsenaars-Schneider model) even at the level of the Hamiltonian. It is of the form
@@
$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{2dCM}}}(x)=-\sum\limits_{j=1}^{N}p_{j}^{2}(kq_{j,x}+c)+\frac{1}{Nc}\Big{(}\sum\limits_{j=1}^{N}p_{j}(kq_{j,x}+c)\Big{)}^{2}+\frac{1}{4}\sum\limits_{j=1}^{N}\frac{k^{4}q_{j,xx}^{2}}{kq_{j,x}+c}+}\\ \displaystyle{+\frac{1}{2}\sum\limits_{i\neq j}^{N}\Big{(}(kq_{i,x}+c)^{2}(kq_{j,x}+c)+(kq_{i,x}+c)(kq_{j,x}+c)^{2}-ck^{2}(q_{i,x}-q_{j,x})^{2}\Big{)}\wp(q_{i}-q_{j})+}\\ \displaystyle{+\frac{k^{3}}{2}\sum\limits_{i\neq j}^{N}\Big{(}q_{i,x}q_{j,xx}-q_{i,xx}q_{j,x}\Big{)}\zeta(q_{i}-q_{j})\,,}\par\end{array}$ (1.7)
where $q_{i,x}=\partial_{x}q_{i}(x)$ and $q_{i,xx}=\partial^{2}_{x}q_{i}(x)$. The upper expression generalizes the finite-dimensional Hamiltonian (1.1). In (1.7) $k\in{\mathcal{R}}$ is a free parameter [^1] One can put $k=1$ but we keep it as it is. Its degree shows the total order of derivative with respect to $x$ in the corresponding term.. When all the fields are independent of $x$ one can put $k=0$, and (1.6) turns into (1.1) with the factor $-2c$. The field analogue of the Calogero-Moser model (1.7) was obtained for 2-body case in [21] and [24] and then extended to $N$-body case in [1]. The canonical Poisson brackets
@@
and the Hamiltonian (1.7) provide equations of motion, which are represented in the form of the zero curvature equation known as the Zakharov-Shabat equation [35, 17]:
In the limit to the finite-dimensional case the term $\partial_{x}{V}(z)$ vanishes, and (1.9) becomes the Lax equation (1.3), where $L(z)$ is the limit of $U(z)$, while $M(z)$ coincides with the limit of $V(z)$ up to some expression commuting with $L(z)$.
The conservation laws in the field case are generated by ${\rm tr}(T^{k}(z,2\pi))$, where $T(z,x)\in{\rm Mat}(N,\mathbb{C})$ is the monodromy matrix
[…]
$\begin{array}[]{c}\displaystyle{\{U_{1}(z,x),U_{2}(w,y)\}=}\\ \\ \displaystyle{\Big{(}-k\partial_{x}{\bf r}_{12}(z,w|x)+[U_{1}(z,x),{\bf r}_{12}(z,w|x)]-[U_{2}(w,y),{\bf r}_{21}(w,z|x)]\Big{)}\delta(x-y)-}\\ \\ \displaystyle{-\Big{(}{\bf r}_{12}(z,w|x)+{\bf r}_{21}(w,z|x)\Big{)}\delta^{\prime}(x-y)\,.}\end{array}$ (1.11)
@@
which was treated in [24] as a modification of the affine Higgs bundle. This establishes gauge equivalence between the 1+1 Calogero-Moser model (1.7) and some special Landau-Lifshitz type model. Examples can be found in [2, 3]. The Landau-Lifshitz type model is described by ultralocal and non-dynamical $r$-matrix of the Belavin-Drinfeld type. We will argue that the gauge transformation (1.12) given by the intertwiner matrix of the IRF-Vertex correspondence provides the same relation between dynamical and non-dynamical $r$-matrices as in the finite-dimensional case. In this way we show that the IRF-Vertex transformation maps the 1+1 Calogero-Moser model into the form with ultralocal $r$-matrix structure.
The paper is organized as follows. In the next Section we recall construction of the classical $r$-matrix structures for the Calogero-Moser model and its spin (and multispin) generalization. In Section 3 we present field analogues of the finite-dimensional constructions from Section 2. In Section 4 the field version of the classical IRF-Vertex relation is discussed.
@@ 2.2 Spin and multispin Calogero-Moser model
$\begin{array}[]{c}\displaystyle{\dot{L}^{\hbox{\tiny{spin}}}(z)+[L^{\hbox{\tiny{spin}}}(z),M^{\hbox{\tiny{spin}}}(z)]=\sum\limits_{i,j=1}^{N}E_{ij}\,(S_{ii}-S_{jj})S_{ij}E_{1}(z)f(z,q_{ij})\,.}\end{array}$ (2.11)
@@
Main statement is that the $U$-matrix (3.9) satisfies the Maillet $r$-matrix structure (1.11) with the $r$-matrix for the finite-dimensional Calogero-Moser model (2.6), where positions of particles $q_{i}$ are replaced with the fields $q_{i}(x)$:
@@
In order to prove this statement we use the standard relations for the delta-function and its derivatives:
@@
$\begin{array}[]{c}\displaystyle{[U_{1}(z,x),{\bf r}_{12}(z,w|x)]-[U_{2}(w,y),{\bf r}_{21}(w,z|x)]=}\\ \\ \displaystyle{+\sum\limits_{i\neq j}^{N}E_{ii}\otimes E_{ij}\,f(w,q_{ij}(x))\alpha_{j}^{2}(x)-\sum\limits_{i\neq j}^{N}E_{ij}\otimes E_{ii}\,f(z,q_{ij}(x))\alpha_{j}^{2}(x)+}\\ \displaystyle{+\sum\limits_{i\neq j}^{N}E_{ij}\otimes E_{jj}\,f(z,q_{ij}(x))\alpha_{j}^{2}(x)-\sum\limits_{i\neq j}^{N}E_{ii}\otimes E_{ji}\,f(w,q_{ji}(x))\alpha_{i}^{2}(x)+}\\ \displaystyle{+\sum\limits_{i\neq j}^{N}E_{ij}\otimes E_{ji}\,f(z-w,q_{ij}(x))(\alpha_{i}^{2}(x)-\alpha_{j}^{2}(x))\,.}\end{array}$ (3.24)
@@
$\begin{array}[]{c}\displaystyle{k\partial_{x}{\bf r}_{12}(z,w|x)=\sum\limits_{i\neq j}^{N}E_{ii}\otimes E_{ji}\,f(w,q_{ji}(x))k(q_{j,x}-q_{i,x})+}\\ \displaystyle{+\sum\limits_{i\neq j}^{N}E_{ij}\otimes E_{ji}\,f(z-w,q_{ij}(x))k(q_{i,x}-q_{j,x})\,.}\end{array}$ (3.25)
@@
$\begin{array}[]{c}\displaystyle{\Big{(}{\bf r}_{12}(z,w|x)+{\bf r}_{21}(w,z|x)\Big{)}\delta^{\prime}(x-y)=\sum\limits_{i=1}^{N}E_{ii}\otimes E_{ii}\Big{(}E_{1}(z)+E_{1}(w)\Big{)}k\delta^{\prime}(x-y)+}\\ \displaystyle{+\sum\limits_{i\neq j}^{N}E_{ij}\otimes E_{jj}\,\phi(z,q_{ij}(x))k\delta^{\prime}(x-y)+\sum\limits_{i\neq j}^{N}E_{ii}\otimes E_{ji}\,\phi(w,q_{ji}(x))k\delta^{\prime}(x-y)\,.}\end{array}$ (3.26)
By summing up all the inputs the statement is proved.
@@
$\begin{array}[]{c}\displaystyle{{\bf r}^{\hbox{\tiny{2dSpin}}}_{12}(z,w)=E_{1}(z-w)\,\sum\limits_{i=1}^{N}E_{ii}\otimes E_{ii}+\sum\limits^{N}_{i\neq j}\phi(z-w,q_{i}(x)-q_{j}(x))\,E_{ij}\otimes E_{ji}\,.}\end{array}$ (3.31)
Then the $U$-matrix (3.28) and $r$-matrix (3.31) satisfy the Maillet ultralocal $r$-matrix structure with unwanted term
@@
This is the field analogue of the finite-dimensional relation (2.19). The non-ultralocal term is absent here since the $r$-matrix (3.31) is skew-symmetric, see (2.21). The proof of the relation (3.32) is simple. Indeed, similarly to the finite-dimensional case $U$-matrix (3.28) and $r$-matrix (3.31) satisfy a direct field generalization of (2.19) with the factor $\delta(x-y)$ for all the terms. But (2.19) does not contain the term $\partial_{x}{\bf r}_{12}(z,w|x)$. Its input in the field case
@@
is added to the unwanted term providing $\mu_{i}$ instead of $\mathcal{S}_{ii}$ in (2.19).
@@ 4 Continuous version of the classical IRF-Vertex relation
$\begin{array}[]{c}\displaystyle{r^{\hbox{\tiny{BD}}}_{12}(z)=E_{1}(z)1_{N}\otimes 1_{N}+\sum\limits_{\begin{subarray}{c}a\in\,\mathbb{Z}_{N}\times\mathbb{Z}_{N}\\ a\neq(0,0)\end{subarray}}T_{a}\otimes T_{-a}\exp(2\pi\imath\,\frac{a_{2}z}{N})\,\phi(z,\frac{a_{1}+a_{2}\tau}{N})\,,}\end{array}$ (4.1)
@@
The $r$-matrix (4.1) is non-dynamical and skew-symmetric [^8] A review of its properties can be found for example in [38]., so that only the commutator terms survive in the Maillet bracket (1.11). The corresponding $U$-matrix is the one of the Sklyanin type, i.e.
$\begin{array}[]{c}\displaystyle{U^{\hbox{\tiny{Skl}}}(z,x)=\sum\limits_{\begin{subarray}{c}a\in\,\mathbb{Z}_{N}\times\mathbb{Z}_{N}\\ a\neq(0,0)\end{subarray}}T_{a}\mathcal{S}_{a}(x)\exp(2\pi\imath\,\frac{a_{2}z}{N})\,\phi(z,\frac{a_{1}+a_{2}\tau}{N})\,.}\end{array}$ (4.4)
In the case $N=2$ it is the $U$-matrix of the Landau-Lifshitz model [30]. More general cases including the multi-pole generalizations of (4.4) were studied in [39, 3].
An existence of a certain gauge transformation relating 1+1 models of the Sklyanin-Landau-Lifshitz type on one hand and the models of 1+1 Calogero-Moser field theory on the other hand was mentioned in [24] and explicit change of variables was derived in [2, 4]. At the same time, a similar relation between dynamical and non-dynamical $r$-matrices is widely known. It is the IRF-Vertex correspondence [6]. The relation between $r$-matrices is given by the intertwining matrix
@@
Notice that the formula (2.26) describing the gauge transformation of $r$-matrix is also valid in the field case [^9] We changed $g$ to $G^{-1}$ in (4.8) with respect to (2.26).
$\begin{array}[]{c}\displaystyle{{\bf r}_{12}(z,w|x)\longrightarrow G_{1}(z,x)G_{2}(w,x)\Big{(}{\bf r}_{12}(z,w|x)-G_{1}^{-1}(z,x)\{G_{1}(z,x),U_{2}(w,x)\}-}\end{array}$ (4.8)
$-\frac{1}{2}\,\Big{[}G_{1}^{-1}(z,x)G_{2}^{-1}(w,x)\{G_{1}(z,x),G_{2}(w,x)\},U_{2}(w,x)\Big{]}\Big{)}G^{-1}_{1}(z,x)G^{-1}_{2}(w,x)\,,$

### v2

@@ 1 Introduction
An integrable field analogue of a finite-dimensional integrable system is some integrable infinite-dimensional generalization of the finite-dimensional model. One possibility is to consider a large $N$ limit. This yields collective field theories of hydrodynamical type. In this paper we use a different approach. A set of coordinates $\varphi_{1},...,\varphi_{n}$ on the phase space of the finite-dimensional model is replaced with a set of field variables $\varphi_{1}(x),...,\varphi_{n}(x)$ depending on $x$ – real valued variable considered as a coordinate on a line or a circle. For definiteness we consider $\varphi_{1}(x),...,\varphi_{n}(x)$ to be periodic functions on a circle: $\varphi_{i}(x)=\varphi_{i}(x+2\pi)$. Dynamics provides also dependence on the time variable $t$, that is we deal with the fields depending on the time variable $t$ and the space variable $x$. Integrable field analogues of the finite-dimensional many-body systems are known beginning from 2d Toda field theory [31]. For example, the 2-body closed Toda model is described by the ordinary differential equation ${\ddot{q}}+\frac{1}{2}\sinh(q)=0$, $q=q(t)$. Its field analogue is the Sinh–Gordon (partial differential) equation $\partial_{t}^{2}q-\partial_{x}^{2}q+\frac{1}{2}\sinh(q)=0$, $q=q(t,x)$. Another example is the Euler top in three-dimensional space: ${\dot{\vec{S}}}={\vec{S}}\times J({\vec{S}})$, where $\vec{S}=\vec{S}(t)=(S_{1}(t),S_{2}(t),S_{3}(t))$ is the angular momentum vector. Its field analogue is given by the Landau-Lifshitz equation:
$\begin{array}[]{c}\displaystyle{\partial_{t}{{\vec{S}}}={\vec{S}}\times\partial^{2}_{x}{\vec{S}}+{\vec{S}}\times J({\vec{S}})\,,\qquad\vec{S}=\vec{S}(t,x)\,,}\end{array}$ (1.6)
where $\vec{S}$ is now the magnetization vector in 1-dimensional model of (ferro)magnet.
Integrability in the 1+1 field case is a subtle question. In this paper by integrability of the field theory we mean a generalization of the finite-dimensional construction based on the Lax equation (1.3) and the classical $r$-matrix structure (1.4). Namely, we assume that the equations of motion are represented in the form of the zero curvature equation known as the Zakharov-Shabat equation [39, 18]:
In the limit to the finite-dimensional case (when all the fields become independent of $x$) the term $\partial_{x}{V}(z)$ vanishes, and (1.7) becomes the Lax equation (1.3), where $L(z)$ is the limit of $U(z)$, while $M(z)$ coincides with the limit of $V(z)$ up to some expression commuting with $L(z)$. The conservation laws in the field case are generated by ${\rm tr}(T^{k}(z,2\pi))$, where $T(z,x)\in{\rm Mat}(N,\mathbb{C})$ is the monodromy matrix
[…]
$\begin{array}[]{c}\displaystyle{\{U_{1}(z,x),U_{2}(w,y)\}=}\\ \\ \displaystyle{\Big{(}-k\partial_{x}{\bf r}_{12}(z,w|x)+[U_{1}(z,x),{\bf r}_{12}(z,w|x)]-[U_{2}(w,y),{\bf r}_{21}(w,z|x)]\Big{)}\delta(x-y)-}\\ \\ \displaystyle{-\Big{(}{\bf r}_{12}(z,w|x)+{\bf r}_{21}(w,z|x)\Big{)}k\delta^{\prime}(x-y)\,.}\end{array}$ (1.9)
In contrast to the Toda model the field generalization of the Calogero-Moser model is highly non-trivial (see also [43] for the field analogue of the Ruijsenaars-Schneider model) even at the level of the Hamiltonian. It is of the form
@@
$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{2dCM}}}(x)=-\sum\limits_{j=1}^{N}p_{j}^{2}(kq_{j,x}+c)+\frac{1}{Nc}\Big{(}\sum\limits_{j=1}^{N}p_{j}(kq_{j,x}+c)\Big{)}^{2}+\frac{1}{4}\sum\limits_{j=1}^{N}\frac{k^{4}q_{j,xx}^{2}}{kq_{j,x}+c}+}\\ \displaystyle{+\frac{1}{2}\sum\limits_{i\neq j}^{N}\Big{(}(kq_{i,x}+c)^{2}(kq_{j,x}+c)+(kq_{i,x}+c)(kq_{j,x}+c)^{2}-ck^{2}(q_{i,x}-q_{j,x})^{2}\Big{)}\wp(q_{i}-q_{j})+}\\ \displaystyle{+\frac{k^{3}}{2}\sum\limits_{i\neq j}^{N}\Big{(}q_{i,x}q_{j,xx}-q_{i,xx}q_{j,x}\Big{)}\zeta(q_{i}-q_{j})\,,}\end{array}$ (1.11)
where $q_{i,x}=\partial_{x}q_{i}(x)$ and $q_{i,xx}=\partial^{2}_{x}q_{i}(x)$, and the set of fields [^1] Similarly to the finite-dimensional case we consider all the fields to be $\mathbb{C}$-valued. is given by $q_{1}(t,x),...,q_{N}(t,x)$ and $p_{1}(t,x),...,p_{N}(t,x)$ with the canonical Poisson brackets
@@
The expression (1.11) for the field theory Hamiltonian generalizes the finite-dimensional Hamiltonian (1.1). In (1.11) $k\in{\mathcal{R}}$ is a free parameter [^2] One can put $k=1$ but we keep it as it is. Its degree shows the total order of derivative with respect to $x$ in the corresponding term.. When all the fields are independent of $x$ one can put $k=0$, and (1.10) turns into (1.1) with the factor $-2c$. The field analogue of the Calogero-Moser model (1.11) was obtained for 2-body case in [23] and [26] and then extended to $N$-body case in [1]. The canonical Poisson brackets (1.12) and the Hamiltonian (1.11) provide equations of motion, which are represented in the form of the zero curvature equation known as the Zakharov-Shabat equation (1.7), see details in [23, 26, 1]. At the same time the classical $r$-matrix structure for the model (1.11) was unknown, and this is one of the purposes of the paper.
Integrable 1+1 field theories are actively studied at the classical and quantum levels due to their deep relation to gauge theories, strings and conformal field theories, existence of soliton solutions and many other remarkable properties [17, 37, 28, 24, 25, 14, 16]. We hope to include into consideration a large class of models obtained as 1+1 field generalizations of the widely known finite-dimensional integrable systems of Calogero-Ruijsenaars family.
@@
$\begin{array}[]{c}\displaystyle{U(z,x)\rightarrow G(z,x)U(z,x)G^{-1}(z,x)+k\partial_{x}G(z,x)G^{-1}(z,x)\,,}\end{array}$ (1.13)
which was treated in [26] as a modification of the affine Higgs bundle. This establishes gauge equivalence between the 1+1 Calogero-Moser model (1.11) and some special Landau-Lifshitz type model. Examples can be found in [2, 3, 4]. The Landau-Lifshitz type model is described by ultralocal and non-dynamical $r$-matrix of the Belavin-Drinfeld type. We will argue that the gauge transformation (1.13) given by the intertwiner matrix of the IRF-Vertex correspondence provides the same relation between dynamical and non-dynamical $r$-matrices as in the finite-dimensional case. In this way we show that the IRF-Vertex transformation maps the 1+1 field Calogero-Moser model into the form with ultralocal $r$-matrix structure (related to Landau-Lifshitz model).
The paper is organized as follows. In the next Section we recall construction of the classical $r$-matrix structures for the Calogero-Moser model and its spin (and multispin) generalization. In Section 3 we present field analogues of the finite-dimensional constructions from Section 2. In Section 4 the field version of the classical IRF-Vertex relation is discussed. Elliptic functions definitions and properties are given in the Appendix A. A brief review of $R$-matrices under consideration and interrelations between them including the IRF-Vertex relations can be found in the Appendix B.
@@ 2.2 Spin and multispin Calogero-Moser model
$\begin{array}[]{c}\displaystyle{\dot{L}^{\hbox{\tiny{spin}}}(z)+[L^{\hbox{\tiny{spin}}}(z),M^{\hbox{\tiny{spin}}}(z)]=-\sum\limits_{i,j=1}^{N}E_{ij}\,(S_{ii}-S_{jj})S_{ij}E_{1}(z)f(z,q_{ij})\,.}\end{array}$ (2.11)
@@
Main statement is as follows.
Theorem 1
The $U$-matrix (3.9) satisfies the Maillet $r$-matrix structure (1.9) with the $r$-matrix for the finite-dimensional Calogero-Moser model (2.6), where positions of particles $q_{i}$ are replaced with the fields $q_{i}(x)$:
@@
Proof:  In order to prove this statement we use the standard relations for the delta-function and its derivatives:
@@
$\begin{array}[]{c}\displaystyle{[U_{1}(z,x),{\bf r}^{\hbox{\tiny{2dCM}}}_{12}(z,w|x)]-[U_{2}(w,y),{\bf r}^{\hbox{\tiny{2dCM}}}_{21}(w,z|x)]=}\\ \\ \displaystyle{+\sum\limits_{i\neq j}^{N}E_{ii}\otimes E_{ij}\,f(w,q_{ij}(x))\alpha_{j}^{2}(x)-\sum\limits_{i\neq j}^{N}E_{ij}\otimes E_{ii}\,f(z,q_{ij}(x))\alpha_{j}^{2}(x)+}\\ \displaystyle{+\sum\limits_{i\neq j}^{N}E_{ij}\otimes E_{jj}\,f(z,q_{ij}(x))\alpha_{j}^{2}(x)-\sum\limits_{i\neq j}^{N}E_{ii}\otimes E_{ji}\,f(w,q_{ji}(x))\alpha_{i}^{2}(x)+}\\ \displaystyle{+\sum\limits_{i\neq j}^{N}E_{ij}\otimes E_{ji}\,f(z-w,q_{ij}(x))(\alpha_{i}^{2}(x)-\alpha_{j}^{2}(x))\,.}\end{array}$ (3.24)
@@
$\begin{array}[]{c}\displaystyle{k\partial_{x}{\bf r}^{\hbox{\tiny{2dCM}}}_{12}(z,w|x)=\sum\limits_{i\neq j}^{N}E_{ii}\otimes E_{ji}\,f(w,q_{ji}(x))k(q_{j,x}-q_{i,x})+}\\ \displaystyle{+\sum\limits_{i\neq j}^{N}E_{ij}\otimes E_{ji}\,f(z-w,q_{ij}(x))k(q_{i,x}-q_{j,x})\,.}\end{array}$ (3.25)
@@
$\begin{array}[]{c}\displaystyle{\Big{(}{\bf r}^{\hbox{\tiny{2dCM}}}_{12}(z,w|x)+{\bf r}^{\hbox{\tiny{2dCM}}}_{21}(w,z|x)\Big{)}k\delta^{\prime}(x-y)=\sum\limits_{i=1}^{N}E_{ii}\otimes E_{ii}\Big{(}E_{1}(z)+E_{1}(w)\Big{)}k\delta^{\prime}(x-y)+}\\ \displaystyle{+\sum\limits_{i\neq j}^{N}E_{ij}\otimes E_{jj}\,\phi(z,q_{ij}(x))k\delta^{\prime}(x-y)+\sum\limits_{i\neq j}^{N}E_{ii}\otimes E_{ji}\,\phi(w,q_{ji}(x))k\delta^{\prime}(x-y)\,.}\end{array}$ (3.26)
[…]
Modified classical $r$-matrix.
In the end of the paragraph consider also the $U$-matrix of the field Calogero-Moser model in the center of mass frame from the very beginning:
$\begin{array}[]{c}\displaystyle{{\bar{U}}^{\hbox{\tiny{2dCM}}}_{ij}(z)=\delta_{ij}\Big{(}p_{i}+{\bar{\alpha}}_{i}^{2}E_{1}(z)-k\frac{{\bar{\alpha}}_{i,x}}{{\bar{\alpha}}_{i}}\Big{)}+(1-\delta_{ij})\phi(z,q_{i}-q_{j}){\bar{\alpha}}_{j}^{2}\,,}\end{array}$ (3.27)
where
$\begin{array}[]{c}\displaystyle{{\bar{\alpha}}_{j}^{2}=k{\bar{q}}_{j,x}+c}\end{array}$ (3.28)
and
$\begin{array}[]{c}\displaystyle{{\bar{q}}_{j}=q_{j}-\frac{1}{N}\sum\limits_{k=1}^{N}q_{k}\,.}\end{array}$ (3.29)
Straightforward calculation shows that in this case the classical $r$-matrix entering the Maillet bracket is of the form:
$\begin{array}[]{c}\displaystyle{{\bar{\bf r}}^{\hbox{\tiny{2dCM}}}_{12}(z,w|x)={\bf r}^{\hbox{\tiny{2dCM}}}_{12}(z,w|x)-{\bar{l}}_{2}(w,x)\,,\qquad{\bar{l}}_{2}(w,x)=1_{N}\otimes{\bar{l}}(w,x)\,,}\end{array}$ (3.30)
and
$\begin{array}[]{c}\displaystyle{{\bar{l}}_{ij}(w,x)=\frac{1}{N}\Big{(}\delta_{ij}E_{1}(w)+(1-\delta_{ij})\phi(w,q_{i}(x)-q_{j}(x))\Big{)}\,.}\end{array}$ (3.31)
@@
$\begin{array}[]{c}\displaystyle{{\bf r}^{\hbox{\tiny{2dSpin}}}_{12}(z,w|x)=E_{1}(z-w)\,\sum\limits_{i=1}^{N}E_{ii}\otimes E_{ii}+\sum\limits^{N}_{i\neq j}\phi(z-w,q_{i}(x)-q_{j}(x))\,E_{ij}\otimes E_{ji}\,.}\end{array}$ (3.36)
Theorem 2
Then the $U$-matrix (3.33) and $r$-matrix (3.36) satisfy the Maillet ultralocal $r$-matrix structure with the unwanted term
@@
Proof: This is the field analogue of the finite-dimensional relation (2.19). The non-ultralocal term is absent here since the $r$-matrix (3.36) is skew-symmetric, see (2.21). The proof of the relation (3.37) is simple. Indeed, similarly to the finite-dimensional case $U$-matrix (3.33) and $r$-matrix (3.36) satisfy a direct field generalization of (2.19) with the factor $\delta(x-y)$ for all the terms. But (2.19) does not contain the term $\partial_{x}{\bf r}_{12}(z,w|x)$. Its input in the field case
@@
[…]
@@ 4 Continuous version of the classical IRF-Vertex relation
$\begin{array}[]{c}\displaystyle{r^{\hbox{\tiny{BD}}}_{12}(z,w)=E_{1}(z-w)\frac{1_{N}\otimes 1_{N}}{N}+\sum\limits_{\begin{subarray}{c}a\in\,\mathbb{Z}_{N}\times\mathbb{Z}_{N}\\ a\neq(0,0)\end{subarray}}\frac{T_{a}\otimes T_{-a}}{N}\exp(2\pi\imath\,\frac{a_{2}(z-w)}{N})\,\phi(z-w,\frac{a_{1}+a_{2}\tau}{N})\,,}\end{array}$ (4.1)
@@
The $r$-matrix (4.1) is non-dynamical, that is it is independent of dynamical variables. In particular, it means that this $r$-matrix is exactly the same in the field case:
$\begin{array}[]{c}\displaystyle{{\bf r}^{\hbox{\tiny{BD}}}_{12}(z,w|x)=r^{\hbox{\tiny{BD}}}_{12}(z,w)}\end{array}$ (4.4)
since
$\begin{array}[]{c}\displaystyle{\partial_{x}{\bf r}^{\hbox{\tiny{BD}}}_{12}(z,w|x)=0\,.}\end{array}$ (4.5)
Moreover, $r$-matrix (4.1) is skew-symmetric [^9] A review of its properties can be found for example in [43].: $r^{\hbox{\tiny{BD}}}_{12}(z,w)=-r^{\hbox{\tiny{BD}}}_{21}(w,z)$, so that only the commutator terms survive in the Maillet bracket (1.9). The corresponding $U$-matrix is the one of the Sklyanin type, i.e.
$\begin{array}[]{c}\displaystyle{U^{\hbox{\tiny{Skl}}}(z,x)=\sum\limits_{\begin{subarray}{c}a\in\,\mathbb{Z}_{N}\times\mathbb{Z}_{N}\\ a\neq(0,0)\end{subarray}}T_{a}\mathcal{S}_{a}(x)\exp(2\pi\imath\,\frac{a_{2}z}{N})\,\phi(z,\frac{a_{1}+a_{2}\tau}{N})\,,}\end{array}$ (4.6)
where $\mathcal{S}_{a}(x)$ are dynamical field variables. From the above arguments it follows that
$\begin{array}[]{c}\displaystyle{\{U^{\hbox{\tiny{Skl}}}_{1}(z,x),U^{\hbox{\tiny{Skl}}}_{2}(w,y)\}=}\\ \\ \displaystyle{=\Big{(}[U^{\hbox{\tiny{Skl}}}_{1}(z,x),{\bf r}^{\hbox{\tiny{BD}}}_{12}(z,w|x)]-[U^{\hbox{\tiny{Skl}}}_{2}(w,y),{\bf r}^{\hbox{\tiny{BD}}}_{21}(w,z|x)]\Big{)}\delta(x-y)=}\\ \\ \displaystyle{=[U^{\hbox{\tiny{Skl}}}_{1}(z,x)+U^{\hbox{\tiny{Skl}}}_{2}(w,x),{\bf r}^{\hbox{\tiny{BD}}}_{12}(z,w|x)]\delta(x-y)\,.}\end{array}$ (4.7)
This relation is equivalent to the Poisson brackets (3.32) for the components of matrix $\mathcal{S}$. In the case $N=2$ the $U$-matrix (4.6) is the one for the Landau-Lifshitz model [34], and three fields $\mathcal{S}_{a}(x)$ are the components of the magnetization vector entering (1.6). More general cases including the multi-pole generalizations and trigonometric and rational degenerations of (4.6) were studied in [44, 3].
An existence of a certain gauge transformation relating 1+1 models of the Sklyanin-Landau-Lifshitz type on one hand and the models of 1+1 Calogero-Moser field theory on the other hand was mentioned in [26] and explicit change of variables was derived in [2, 4]. At the same time, a similar relation between dynamical and non-dynamical $r$-matrices is widely known. It is the IRF-Vertex correspondence [7]. See a brief review in the Appendix B. The relation between $r$-matrices is written in terms of intertwining matrix
@@
Notice that the formula (2.26) describing the gauge transformation of $r$-matrix is also valid in the field case. Consider a gauge transformation with ”ultra-local brackets”, that is suppose that the Poisson brackets for a matrix $G(z,x)\in{\rm Mat}(N,\mathbb{C})$ of gauge transformation have the form
$\begin{array}[]{c}\displaystyle{\{G_{1}(z,x),U_{2}(w,y)\}=\langle\{G_{1},U_{2}\}\rangle(z,w,x)\delta(x-y)\,,}\\ \\ \displaystyle{\{G_{1}(z,x),G_{2}(w,y)\}=\langle\{G_{1},G_{2}\}\rangle(z,w,x)\delta(x-y)\,,}\end{array}$ (4.11)
where the expressions $\langle\{G_{1},U_{2}\}\rangle(z,w,x)$ and $\langle\{G_{1},G_{2}\}\rangle(z,w,x)$ are some ${\rm Mat}(N,\mathbb{C})^{\otimes 2}$-valued functions of the field variables obtained by direct calculation from a given Poisson structure.
Lemma 1
Suppose $U\in{\rm Mat}(N,\mathbb{C})$ satisfies the Maillet bracket (1.9) with some $r$-matrix ${\bf r}_{12}(z,w,x)$. Consider a gauge transformation
with a matrix of gauge transformation satisfying the property of ultra-local brackets (4.11). Then the gauge transformed $U$-matrix $U^{\prime}(z,x)=G(z,x)U(z,x)G^{-1}(z,x)+k\partial_{x}G(z,x)G^{-1}(z,x)$ again satisfies the Maillet bracket (1.9) with the gauge transformed $r$-matrix
$\begin{array}[]{c}\displaystyle{{\bf r}_{12}(z,w|x)\longrightarrow G_{1}(z,x)G_{2}(w,x)\Big{(}{\bf r}_{12}(z,w|x)-G_{1}^{-1}(z,x)\langle\{G_{1},U_{2}\}\rangle(z,w,x)-}\end{array}$ (4.13)
$-\frac{1}{2}\,\Big{[}G_{1}^{-1}(z,x)G_{2}^{-1}(w,x)\langle\{G_{1},G_{2}\}\rangle(z,w,x),U_{2}(w,x)\Big{]}\Big{)}G^{-1}_{1}(z,x)G^{-1}_{2}(w,x)\,.$